OpenStudy (anonymous):

Can someone show me how to factor this step by step: 8x^2+18x+9

OpenStudy (anonymous):

For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (72) that add up to b (18).In this problem (3)/(2)*(3)/(4)=(9)/(8) (which is (c)/(a)) and (3)/(2)+(3)/(4)=(18)/(8) (which is ((b)/(a)) , so insert (3)/(2) as the right hand term of one factor and (3)/(4) as the right-hand term of the other factor. (x+(3)/(2))(x+(3)/(4))

OpenStudy (anonymous):

Remove the fraction by multiplying the first term of the factor by the denominator of the second term. (2x+3)(4x+3)

OpenStudy (anonymous):

thankx..

OpenStudy (anonymous):

Hope it made sense.

OpenStudy (campbell_st):

here is a really simply method for a quadratic ax^2 + bx + c = 0 multiply a and c in this question its 8 x 9 = 72 find the factors of 72 that add to 18.... 12 + 6 so then you have  remove common factors from each binomial  delete the common factor from the numerator and denominator... 8 leaves (2x + 3)(4x + 3) = 0 just solve now for x.... this method works all the time...